OpenStudy (anonymous):

Four employees at a crayon factory are working on four separate machines. Cynthia's machine produces 100 less than twice the number of crayons as Zane's machine every hour. Louisa's machine produces 50 crayons more than Zane's machine every hour. Evan's machine produces 20 crayons less than Cynthia's machine every hour. How many crayons must Zane's machine produce every hour in order for Cynthia and Zane output to equal Louisa and Evan's output?

OpenStudy (yttrium):

Let x be Zane's machine. Cynthia then will have 2x - 100 Louisa will have x + 50 Evan will have x - 20 Just equate Zane+Cynthia and Louisa+Evan's machine's output and then solve for x. Can you do it?

OpenStudy (anonymous):

i'm not sure very bad with word problems but if I did it correctly is x=-70 if not what step is wrong

OpenStudy (yttrium):

Did you arrive at this way? x+2x-100 = x+50+x-20 3x-100 = 2x +30 Therefore, what is x?

OpenStudy (anonymous):

yes

OpenStudy (yttrium):

What is your x, then? Maybe you just made some error on calculations. It's not -70 btw. :)

OpenStudy (anonymous):

ok is it 5x70

OpenStudy (anonymous):

sorry 5x-70

OpenStudy (anonymous):

I am confused can you walk me through this one I will mimic the formula for my similar problems

OpenStudy (yttrium):

You just apply the law of axioms or transpositions. We are now with 3x-100 = 2x+30, right? Using tranposition (you need to change the sign if you transpose), see the process below 3x-100 - 2x - 30 = 0 Therefore, x - 130 = 0 Therefore, x is what?

OpenStudy (anonymous):

I think I got it x=130